Situation:



Suppose you were one of the engineers of the said project and your job was to renovate or improve the walkway, patio, and driveway. After your ocular inspection, you noticed that a rectangular floor measuring (10m by 14m) needed to be fixed. Likewise, your plan is to put brick paves to ensure that the walkway is strong and durable.




Questions:



1. Each piece of brick is square with an edge of 50 cm. How many pieces of brick paves will be needed to cover the rectangular floor that needs fixing?



2. If one bag of adhesive cement for brick paves can cover 10 sq. m, how many bags of adhesive cement will be needed?



3. Make a model to illustrate the situation with appropriate mathematical solutions.

1
Expert's answer
2021-11-16T06:53:02-0500

"\text{The area of the floor is given by $14 . 10 = 140cm^2$, also the area of }\\
\text{each brick is given by $0.5\cdot 0.5=0.25cm^2$, the number of bricks is}\\\text{given by $\frac{140}{0.25}=560$}\\
\text{Also if one bag of cement covers $10m^2$, then $\frac{140}{10} = 14 $ bags }\\ \text{covers the whole floor }" The third part of the question is not clear as it doesn't seem to correlate with question